OpenStudy (anonymous):

What is the simplified form of 4x plus 2 over x plus 5 plus the fraction 3x plus 1 over x plus 5? (I know that the question is written in a weird context, but that is exactly how it is being asked.)

jimthompson5910 (jim_thompson5910):

so you have \[\large \frac{4x+2}{x+5} + \frac{3x+1}{x+5}\] ???

OpenStudy (anonymous):

Yes, that is how it writes out to be. Would you like me to post the multiple choice answers for help?

jimthompson5910 (jim_thompson5910):

no I know how to do this

OpenStudy (anonymous):

Ok, that's what I thought.

jimthompson5910 (jim_thompson5910):

The denominators are the same. So this means you can add the numerators and place that result over the same common denominator.

jimthompson5910 (jim_thompson5910):

The numerators are 4x+2 and 3x+1 add them: 4x+2+3x+1 = ???

OpenStudy (anonymous):

hold on...

jimthompson5910 (jim_thompson5910):

alright

OpenStudy (anonymous):

7x +3

jimthompson5910 (jim_thompson5910):

you then place that over the LCD, which is x+5

jimthompson5910 (jim_thompson5910):

so this is what you're doing \[\large \frac{4x+2}{x+5} + \frac{3x+1}{x+5}\] \[\large \frac{(4x+2)+(3x+1)}{x+5}\] \[\large \frac{4x+2+3x+1}{x+5}\] \[\large \frac{(4x+3x)+(2+1)}{x+5}\] \[\large \frac{7x+3}{x+5}\]

jimthompson5910 (jim_thompson5910):

again, I must emphasize this: you can only add the numerators ONLY if the denominators are the same

OpenStudy (anonymous):

Fantastic explanation! Thank you very much. Mathematics do not come easy to me, I am more of a "right-brained" person. Although that is no excuse. :P Thanks again @jim_thompson5910

jimthompson5910 (jim_thompson5910):

I'm sure with a lot of hard work, it will click eventually. So don't give up.

jimthompson5910 (jim_thompson5910):

Glad to be of help

OpenStudy (anonymous):

:)